# PepCompass: Navigating peptide embedding spaces using Riemannian Geometry

### v1

@@ 1 Introduction
Efficient exploration of peptide space is notoriously difficult. At the global level, there are more than $3.3\times 10^{32}$ combinatorially possible amino acid sequences of length at most 25. At the local level, each peptide of length 25 has nearly 1000 neighbors within an edit radius of one. Moreover, only a small fraction of amino acid sequences correspond to antimicrobial peptides (AMPs), which have high activity against bacteria Szymczak2023AIAMPDiscovery ; Szymczak2025AIDrivenAMP . This extreme combinatorial complexity renders AMP discovery by brute-force exploration intractable. To address this challenge, we turn to one of the most prolific inventions of humankind: maps.
Since the dawn of civilization, maps have provided a structured way to support both local and global navigation, driving scientific discovery. In modern machine learning, latent-space generative models such as VAEs, GANs, WAEs, and normalizing flows (BondTaylor2021ComparativeReview, ) enable building continuous latent representations—maps—of peptides. Such maps have already facilitated the discovery of promising new AMPs (Szymczak2023HydrAMP, ; VanOort2021AMPGANv2, ; Wang2022LSSAMP, ; Das2020CLaSS, ). The standard workflow assumes that once the model is trained, its latent space together with the decoder properly models a set of valid, synthetizable peptides. The latent space is typically chosen to be $\mathbb{R}^{d}$ with a flat Euclidean metric, enabling direct application of existing exploration and optimization algorithms. However, such flat representations suffer from a significant flaw: they ignore the differential geometry induced by the decoder, leading to distortions in distances.
Typical approaches attempting to circumvent this problem assume the manifold hypothesis (Bengio2012RepresentationLearning, ) and use the pullback metric (arvanitidis2018latent, ). However, recent work (LoaizaGanem2024ManifoldSurvey, ; brown2022uomh, ; wang2024cw, ) has shown that the manifold hypothesis does not withstand empirical scrutiny for image data, where sets of images are better modeled as unions or CW-complexes of manifolds with varying low dimensionality (we refer to Appendix A for Related Work). We show that peptide spaces suffer from a similar issue and introduce decoder-derived Union of $\kappa$-Stable Riemannian Manifolds $\mathbb{M}^{\kappa}$, which captures both the complex structure of peptide space and its local geometry, with computational stability controlled by a parameter $\kappa$. Intuitively, we cut the globally distorted map into a set of charts that enable efficient and distortion-free exploration and optimization.
@@
For local search on a single manifold from the family $\mathbb{M}^{\kappa}$, we designed two geometry-informed approaches: Second-Order Riemannian Brownian Efficient Sampling (SORBES) and Mutation Enumeration in Tangent Space (MUTANG). SORBES is a provably convergent, second-order approximation of the Riemannian Brownian motion (schwarz2022randomwalk, ; herrmann2023subriemannian, ), serving as a Riemannian analogue of local Gaussian search. MUTANG addresses the discrete nature of peptide space by reinterpreting the local tangent space not as continuous vectors but as discrete mutations, directly corresponding to amino-acid substitutions. This reinterpretation provides both interpretability and efficiency, enabling enumeration of a given peptide’s neighbours. We further combine SORBES and MUTANG into an iterative Local Enumeration procedure that densely populates the neighbourhood of a given peptide with valid, diverse neighbours. Finally, integrating this enumeration with a Bayesian optimization scheme yields an efficient Local Enumeration Bayesian Optimization procedure (LE-BO; see Figure 1B,C).
@@ 2.1 Background
A common approach is to equip $\mathcal{Z}$ with the standard Euclidean inner product $\langle\cdot,\cdot\rangle_{\mathbb{R}^{d}}$, and use the associated Euclidean distance as a base for exploration. However, this ignores the geometry induced by the decoder, which, under the manifold hypothesis, can be accounted for using the pullback metric (Bengio2012RepresentationLearning, ; arvanitidis2018latent, ).
@@
Previous work demonstrates that the manifold hypothesis often fails for complex data such as images (LoaizaGanem2024ManifoldSurvey, ; brown2022uomh, ; wang2024cw, ), and the data is better represented as unions of local manifolds of varying dimension, typically lower than that of the latent space. However, the previous methods defined the submanifolds based on pre-specified datasets and could not generalize to new data.
@@ 2.2 Union of $\kappa$-Stable Riemannian Manifolds
Assuming that the manifold hypothesis is indeed violated and that a given generative model has learned to faithfully capture the lower-dimensional structure in the data, it should be reflected in the decoder having rank strictly smaller than the latent dimensionality. We verified this phenomenon for antimicrobial peptides data in two state-of-the-art latent generative models (Das2018PepCVAE, ; Szymczak2023HydrAMP, ) (see Appendix B). This implies that the $G_{Dec}$ is not of full rank, and consequently the pair $(\mathcal{Z},G_{\operatorname{Dec}})$ does not constitute a Riemannian manifold.
@@ 2.3 SORBES - Second-Order Riemannian Brownian efficient Sampling
For the proof, see Appendix D. Intuitively, in the small–step limit, our algorithm converges to Riemannian Brownian motion on $M_{z}^{\kappa}(\alpha)$, stopped at the boundary, with the deviation probability decaying exponentially in the inverse time horizon. Theorem 1 extends the main convergence result of schwarz2022randomwalk to possibly non-compact manifolds. Importantly, schwarz2022randomwalk showed that achieving this convergence requires a second-order correction term (capturing the effect of Christoffel symbols), rather than the commonly used naive first-order update. This motivates our use of the second-order scheme in Algorithm 1. In Appendix E (Algorithm 4), we describe SORBES-SE (Stable/Efficient), an implementation of SORBES that approximates the second-order correction using finite differences. It employs an adaptive step size $\epsilon$ that adjusts to the local curvature of the space, while still preserving the convergence guarantees.
@@ 2.6 LE-BO - Local Enumeration Bayesian Optimization
Finally, we integrate our LocalEnumeration algorithm into a Bayesian optimization (Garnett2023BayesianOptimization, ) framework for peptide design, which we term Local Enumeration Bayesian Optimization (LE-BO). Instead of performing costly optimization of the acquisition function in the latent space, which typically relies on Euclidean-distance kernels and ignores both the latent geometry and the discrete nature of peptides, we use surrogate Gaussian process models (Rasmussen2006GPML, ), defined directly in the peptide space. The acquisition function is optimized by locally enumerating peptides in the vicinity of the most promising candidates and then selecting the peptide that maximizes the acquisition value. To further encourage exploration and increase the diversity of discovered candidates, we employ the ROBOT scheme (Maus2023ROBOT, ), which promotes searching across a broader set of promising regions. Details of LE-BO are presented in Algorithm 3.
@@ 3.2 LE-BO evaluation
We next evaluated our LE-BO algorithm on a black-box peptide optimization task with a budget of 1400 evaluations. Optimization was initialized from four seed peptides derived using PoGS for $\lambda~=0.01$ and $\mu=0.1$ (Sec. 2.7): KY14, KF16, KK16, FL14, as well as two previously described AMPs (mammuthusin-3, hydrodamin-2(Wan2024DeepLearningEnabled, ). The optimization procedure was repeated 10 times, and for each run the best value achieved across all optimization steps was recorded and reported in Table 2. LE-BO hyperparameters are described in Appendix I.
@@
As the optimized black-box function, we used the APEX MIC regressor (Wan2024DeepLearningEnabled, ), with the objective of minimization. We minimize the average $\log_{2}$ MIC across three E. coli strains, reporting the mean over the best results from 10 repeated optimization runs (Table 2).
@@ 3.3 Wet-lab validation
To validate the computational predictions from PoGS and LE-BO, we conducted comprehensive in vitro antimicrobial testing of the generated peptides. A total of 29 novel peptides were experimentally evaluated: 4 seed peptides discovered through PoGS bi-prototype geodesics with property-aware potentials, and 25 analogs derived from these 4 seeds through LE-BO optimization for E. coli activity. These peptides were tested against a panel of 19 bacterial strains, including 8 multidrug-resistant (MDR) isolates (Appendix J), to assess both broad-spectrum activity and efficacy against clinically relevant resistant pathogens. We compared the success rate of PepCompass to previous methods that were also validated experimentally (HydrAMP (Szymczak2023HydrAMP, ), AMP-Diffusion (torres2025ampdiff, ), CLaSS (Das2020CLaSS, ), Joker (Porto2018Joker, )). To this end, for each activity threshold, we computed the fraction of tested peptides that were active against at least one bacterial strain with this activity thershold.
@@ Appendix C Construction of $\kappa$-Stable Manifolds
$\bar{W}^{\kappa}_{z}\cap(W_{z}^{\kappa})^{\perp}\subset\operatorname{span}\{V(z)_{:,k_{z}^{\kappa}},\dots,V(z)_{:,d}\}.$
@@
$\langle v,v\rangle^{\text{Dec}}_{z}$ $=a_{k_{z}^{\kappa}}^{2}\sigma_{k_{z}^{\kappa}}^{2}\|V(z)_{:,k_{z}^{\kappa}}\|^{2}+\dots+a_{d-1}^{2}\sigma_{d-1}^{2}\|V(z)_{:,d}\|^{2}$ (11)
@@ Appendix D Proof of Theorem 1
Let $\overline{M^{\kappa}_{z}(\alpha)}$ be the spherical-cap compactification of $M^{\kappa}_{z}(\alpha)$. schwarz2022randomwalk [13] showed that the non-stopped version of Algorithm 1, extended to $\overline{M^{\kappa}_{z}(\alpha)}$, produces a process $\overline{Z^{\epsilon}}$ that converges to $B$ in the Skorokhod topology on $\overline{M^{\kappa}_{z}(\alpha)}$.
@@ Appendix E SORBES implementation details
We adopt the efficient implementation of schwarz2022randomwalk [13], which exploits a thin SVD of the decoder Jacobian to orthogonalize tangent directions.

### v2

@@ 1 Introduction
Efficient exploration of peptide space is notoriously difficult. At the global level, there are more than $3.3\times 10^{32}$ combinatorially possible amino acid sequences of length at most 25. At the local level, each peptide of length 25 has nearly 1000 neighbors within an edit radius of one. Moreover, only a small fraction of amino acid sequences correspond to antimicrobial peptides (AMPs), which have high activity against bacteria [1, 2]. This extreme combinatorial complexity renders AMP discovery by brute-force exploration intractable. To address this challenge, we turn to one of the most prolific inventions of humankind: maps.
Since the dawn of civilization, maps have provided a structured way to support both local and global navigation, driving scientific discovery. In modern machine learning, latent-space generative models such as VAEs, GANs, WAEs, and normalizing flows [3] enable building continuous latent representations—maps—of peptides. Such maps have already facilitated the discovery of promising new AMPs [4, 5, 6, 7]. The standard workflow assumes that once the model is trained, its latent space together with the decoder properly models a set of valid, synthetizable peptides. The latent space is typically chosen to be $\mathbb{R}^{d}$ with a flat Euclidean metric, enabling direct application of existing exploration and optimization algorithms. However, such flat representations suffer from a significant flaw: they ignore the differential geometry induced by the decoder, leading to distortions in distances.
Typical approaches attempting to circumvent this problem assume the manifold hypothesis [8] and use the pullback metric [9]. However, recent work [10, 11, 12] has shown that the manifold hypothesis does not withstand empirical scrutiny for image data, where sets of images are better modeled as unions or CW-complexes of manifolds with varying low dimensionality (we refer to Appendix A for Related Work). We show that peptide spaces suffer from a similar issue and introduce decoder-derived Union of $\kappa$-Stable Riemannian Manifolds $\mathbb{M}^{\kappa}$, which captures both the complex structure of peptide space and its local geometry, with computational stability controlled by a parameter $\kappa$. Intuitively, we cut the globally distorted map into a set of charts that enable efficient and distortion-free exploration and optimization.
@@
For local search on a single manifold from the family $\mathbb{M}^{\kappa}$, we designed two geometry-informed approaches: Second-Order Riemannian Brownian Efficient Sampling (SORBES) and Mutation Enumeration in Tangent Space (MUTANG). SORBES is a provably convergent, second-order approximation of the Riemannian Brownian motion [13, 14], serving as a Riemannian analogue of local Gaussian search. MUTANG addresses the discrete nature of peptide space by reinterpreting the local tangent space not as continuous vectors but as discrete mutations, directly corresponding to amino-acid substitutions. This reinterpretation provides both interpretability and efficiency, enabling enumeration of a given peptide’s neighbours. We further combine SORBES and MUTANG into an iterative Local Enumeration procedure that densely populates the neighbourhood of a given peptide with valid, diverse neighbours. Finally, integrating this enumeration with a Bayesian optimization scheme yields an efficient Local Enumeration Bayesian Optimization procedure (LE-BO; see Figure 1B,C).
@@ 2.1 Background
A common approach is to equip $\mathcal{Z}$ with the standard Euclidean inner product $\langle\cdot,\cdot\rangle_{\mathbb{R}^{d}}$, and use the associated Euclidean distance as a base for exploration. However, this ignores the geometry induced by the decoder, which, under the manifold hypothesis, can be accounted for using the pullback metric [8, 9].
@@
Previous work demonstrates that the manifold hypothesis often fails for complex data such as images [10, 11, 12], and the data is better represented as unions of local manifolds of varying dimension, typically lower than that of the latent space. However, the previous methods defined the submanifolds based on pre-specified datasets and could not generalize to new data.
@@ 2.2 Union of $\kappa$-Stable Riemannian Manifolds
Assuming that the manifold hypothesis is indeed violated and that a given generative model has learned to faithfully capture the lower-dimensional structure in the data, it should be reflected in the decoder having rank strictly smaller than the latent dimensionality. We verified this phenomenon for antimicrobial peptides data in two state-of-the-art latent generative models [16, 4] (see Appendix B). This implies that the $G_{Dec}$ is not of full rank, and consequently the pair $(\mathcal{Z},G_{\operatorname{Dec}})$ does not constitute a Riemannian manifold.
@@ 2.3 SORBES - Second-Order Riemannian Brownian efficient Sampling
For the proof, see Appendix D. Intuitively, in the small–step limit, our algorithm converges to Riemannian Brownian motion on $M_{z}^{\kappa}(\alpha)$, stopped at the boundary, with the deviation probability decaying exponentially in the inverse time horizon. Theorem 1 extends the main convergence result of Schwarz et al. [13] to possibly non-compact manifolds. Importantly, Schwarz et al. [13] showed that achieving this convergence requires a second-order correction term (capturing the effect of Christoffel symbols), rather than the commonly used naive first-order update. This motivates our use of the second-order scheme in Algorithm 1. In Appendix E (Algorithm 4), we describe SORBES-SE (Stable/Efficient), an implementation of SORBES that approximates the second-order correction using finite differences. It employs an adaptive step size $\epsilon$ that adjusts to the local curvature of the space, while still preserving the convergence guarantees.
@@ 2.6 LE-BO - Local Enumeration Bayesian Optimization
Finally, we integrate our LocalEnumeration algorithm into a Bayesian optimization [17] framework for peptide design, which we term Local Enumeration Bayesian Optimization (LE-BO). Instead of performing costly optimization of the acquisition function in the latent space, which typically relies on Euclidean-distance kernels and ignores both the latent geometry and the discrete nature of peptides, we use surrogate Gaussian process models [18], defined directly in the peptide space. The acquisition function is optimized by locally enumerating peptides in the vicinity of the most promising candidates and then selecting the peptide that maximizes the acquisition value. To further encourage exploration and increase the diversity of discovered candidates, we employ the ROBOT scheme [19], which promotes searching across a broader set of promising regions. Details of LE-BO are presented in Algorithm 3.
@@ 3.2 LE-BO evaluation
We next evaluated our LE-BO algorithm on a black-box peptide optimization task with a budget of 1400 evaluations. Optimization was initialized from four seed peptides derived using PoGS for $\lambda~=0.01$ and $\mu=0.1$ (Sec. 2.7): KY14, KF16, KK16, FL14, as well as two previously described AMPs (mammuthusin-3, hydrodamin-2[20]. The optimization procedure was repeated 10 times, and for each run the best value achieved across all optimization steps was recorded and reported in Table 2. LE-BO hyperparameters are described in Appendix I.
@@
As the optimized black-box function, we used the APEX MIC regressor [20], with the objective of minimization. We minimize the average $\log_{2}$ MIC across three E. coli strains, reporting the mean over the best results from 10 repeated optimization runs (Table 2).
@@ 3.3 Wet-lab validation
To validate the computational predictions from PoGS and LE-BO, we conducted comprehensive in vitro antimicrobial testing of the generated peptides. A total of 29 novel peptides were experimentally evaluated: 4 seed peptides discovered through PoGS bi-prototype geodesics with property-aware potentials, and 25 analogs derived from these 4 seeds through LE-BO optimization for E. coli activity. These peptides were tested against a panel of 19 bacterial strains, including 8 multidrug-resistant (MDR) isolates (Appendix J), to assess both broad-spectrum activity and efficacy against clinically relevant resistant pathogens. We compared the success rate of PepCompass to previous methods that were also validated experimentally (HydrAMP [4], AMP-Diffusion [35], CLaSS [7], Joker [30]). To this end, for each activity threshold, we computed the fraction of tested peptides that were active against at least one bacterial strain with this activity thershold.
@@ Appendix C Construction of $\kappa$-Stable Manifolds
$\bar{W}^{\kappa}_{z}\cap(W_{z}^{\kappa})^{\perp}\subset\operatorname{span}\{V(z)_{:,k_{z}^{\kappa}},\dots,V(z)_{:,d-1}\}.$
@@
$\langle v,v\rangle^{\text{Dec}}_{z}$ $=a_{k_{z}^{\kappa}}^{2}\sigma_{k_{z}^{\kappa}}^{2}\|V(z)_{:,k_{z}^{\kappa}}\|^{2}+\dots+a_{d-1}^{2}\sigma_{d-1}^{2}\|V(z)_{:,d-1}\|^{2}$ (11)
@@ Appendix D Proof of Theorem 1
Let $\overline{M^{\kappa}_{z}(\alpha)}$ be the spherical-cap compactification of $M^{\kappa}_{z}(\alpha)$. Schwarz et al. [13] showed that the non-stopped version of Algorithm 1, extended to $\overline{M^{\kappa}_{z}(\alpha)}$, produces a process $\overline{Z^{\epsilon}}$ that converges to $B$ in the Skorokhod topology on $\overline{M^{\kappa}_{z}(\alpha)}$.
@@ Appendix E SORBES implementation details
We adopt the efficient implementation of Schwarz et al. [13], which exploits a thin SVD of the decoder Jacobian to orthogonalize tangent directions.